Draw eight cards from a standard deck of cards with replacement.
What is the probability that of these first eight cards that five are spades?
I come up with =====.001981?

Answer by stanbon(75887)

Draw eight cards from a standard deck of cards with replacement.
What is the probability that of these first eight cards that five are spades?
------
Binomial Problem with n = 8 and p(spade) = 1/4 ; P (not spade = 3/4)
P(x = 5) = 8C5(1/4)^5*(3/4)^3 = 0.02307
==================================
Cheers,
Stan H.
============